the I.Sc lecture theatre of a college is 40 ft wide and has a door at a corner. A teacher enters at 12.00 noon through the door and makes 10 rounds along the 40 ft wall back and fortg during the period and finally leaves the classroom at 12.50 p.m. through the same door. compute his average speed and average velocity.

Grade:11

3 Answers

Arun
6 years ago

Total time for which the teacher moves within the room, t = 50 min = 3000 s

In one round the teacher covers =  40 + 40 = 80 ft

So, the total distance covered by the teacher in 10 rounds is = 10 × 80 = 800 ft

Now, average speed = total distance/total time = 800/3000 = (4/15) ft/s

Now, for average velocity, the net displacement of the teacher is zero. Since, he leaves the room through the same door through which he entered, the average velocity will be zero.

Mahima churi
5 years ago
In one round teacher cover 40 +40 =80 ft 
Doin 10 rounds it will cover 80×10 =800 
Avg speed =800÷50=16ft/min
Since there was no displacmdisp therefore avg velocity is zero
ananya
5 years ago
total speed travelled in 50mon = 800 ft.
average speed = 800/50 ft/min
                        = 16 ft/min.
At 12:00 noon he is at the door and at 12:50 he is again at the same door.
therefore the displacementduring the 50 mins is 0.
Average velocity is also equal to 0